# Are EU low-carbon structural funds efficient in reducing emissions?

### v1

@@ 4.2 Effect in the emissions cycle
| GDP | 0.458∗∗∗ | 0.519∗∗∗ | 0.538∗∗∗ | 0.463∗∗∗ | ||
| (0.028) | (0.035) | (0.028) | (0.035) | |||
| EPS | 0.025∗∗∗ | 0.030∗∗∗ | 0.017∗∗∗ | 0.016∗∗∗ | ||
| (0.004) | (0.004) | (0.004) | (0.004) | |||
| Political Ideology | 0.003 | 0.002 | 0.001 | 0.0001 | ||
| (0.003) | (0.003) | (0.003) | (0.003) | |||
| GVAA | 0.011 | 0.006 | ||||
| (0.008) | (0.008) | |||||
| GVAB-E | 0.112∗∗∗ | 0.082∗∗∗ | ||||
| (0.016) | (0.016) | |||||
| GVAF | 0.086∗∗∗ | 0.030∗∗ | ||||
| (0.013) | (0.013) | |||||
| GVAG-J | 0.058∗∗ | 0.199∗∗∗ | ||||
| (0.025) | (0.025) | |||||
| GVAK-N | 0.069∗∗ | 0.143∗∗∗ | ||||
| (0.033) | (0.033) | |||||
| GVAO-U | 0.187∗∗∗ | 0.0003 | ||||
| (0.037) | (0.037) | |||||
| Observations | 3,872 | 2,996 | 2,996 | 3,872 | 2,996 | 2,996 |
| R2 | 0.248 | 0.295 | 0.302 | 0.166 | 0.150 | 0.136 |
| Adjusted R2 | 0.193 | 0.235 | 0.241 | 0.105 | 0.078 | 0.060 |
For transition regions, the estimated effect is relatively clearer; nonetheless, determining the net effect related to each program has opposite signs. The interaction with expenditures in the 2007-2013 programme shows a positive and significant effect (except in models 1 and 4) even after discounting the effect of the reference group. The net effect of expenditure in this period suggests an increase in the intensity of emission cycles, both per capita and concerning GDP. In contrast, the effect associated with expenditure in the 2014-2020 programme is consistently negative and significant (except in models 1 and 4). Then, considering the effect of the reference group, the net effect remains negative and comparatively much stronger. Although the estimated effects for both periods are opposing, the expected effect of these expenditures is to reduce the intensity of emission cycles. For less developed regions, the effect of expenditures associated with emission cycles is negative in almost all specifications for both periods, although not significant. Thus, considering the estimated effects in the reference group, the effect of expenditures on emission cycles in less developed regions is not significant.
Regarding the other control variables, the GDP cycle has a strong positive impact on the per capita emission cycle and a negative impact on the emission intensity cycle, indicating that economic growth cycles are associated with higher per capita emissions but with a lower intensity of emissions intensity. The country-level EPS dummy shows a significant negative effect on emissions, both per capita and intensity, suggesting that stricter environmental policies are associated with lower emissions in their cycles. The region’s weighted average political ideology does not show a significant impact on emission cycles, indicating that local political outcomes, whether left or right, may not have a direct relationship with the cyclical variability of emissions.
Finally, the sectoral gross value added coefficients indicate that different sectors have varying impacts on emission cycles. For example, the manufacturing sector (B-E) has a significant positive impact on per capita emissions and a negative impact on emission intensity. Therefore, most industry sectors have a significant influence on the cyclical variations in emissions.
4.3 Effect by type of gases
We break down emissions by specific gases, considering CO2, CH4, and N2O, to assess low-carbon investments’ impacts, using the same methodology. Detailed numerical result tables can be found in the Appendix (Tables A4 to A9), allowing us to concentrate on qualitative insights.
Disaggregated GHG emissions data at the regional level allow for a better understanding of the heterogeneous impacts of the investments and can help identify specific expenditure gaps to improve policy responses. While CO2 contributes significantly to GHG emissions and changes in CO2 emissions primarily explain the impact of low-carbon funding, it is essential to consider a broader range of gases, which have different sources and economic activity patterns. In addition, decomposing GHG emissions into leading pollutant gases is relevant as European regions vary significantly in emission types, which are linked to their socioeconomic composition (Crippa et al., 2023). CO2 results from burning fossil fuels, chemical processes like cement production, and natural emissions from trees and biological materials. CH4 is emitted during fossil fuel extraction and transportation and significantly from agricultural activities such as livestock and organic waste decomposition in landfills. N2O is produced by agricultural and industrial activities, fossil fuel combustion, and wastewater treatment.
The fixed effects estimations results on CO2, CH4, and N2O emissions are presented on Tables A4 to A6 in the Appendix.
For CO2 emissions, low-carbon expenditures for the 2007-2013 and 2014-2020 programmes show significant negative impacts on emissions per capita and relative to GDP in developed and transition regions, indicating effective reductions. However, in less developed regions, these expenditures result in positive significant impacts, suggesting an increase in emissions or less effective reductions. These results are similar to those observed for overall GHG emissions (Table 2).
For CH4 emissions, low-carbon expenditures for the 2007-2013 programme generally show negative but non-significant impacts across most specifications (including emissions per capita and relative to GDP). For the 2014-2020 programme, there are significant negative impacts in transition regions but positive or non-significant impacts in less developed regions, indicating inconsistent effectiveness. Regarding N2O emissions, low-carbon expenditures for both the 2007-2013 and 2014-2020 programmes have negative impacts but are not significant in developed regions. In transition regions, the impacts are negative and significant, while in less developed regions, the impacts are positive and significant, indicating less effective reductions or increases in emissions.
These results show that low-carbon expenditures have varying impacts depending on the type of gas and the level of regional development. In developed and transition regions, these investments are generally effective in reducing CO2 emissions but exhibit mixed and often non-significant results for CH4 and N2O. In less developed regions, low-carbon expenditures frequently result in increased emissions or less effective reductions across all gases.
Tables A7 to A9 in the Appendix present the fixed effect estimation results on the cyclical components of CO2, CH4, and N2O emissions, respectively. For CO2 emissions, low-carbon expenditures for the 2007-2013 and 2014-2020 programmes show generally non-significant impacts on the cyclical component of emissions per capita and relative to GDP. There are mixed results in transition regions, with significant positive impacts for 2007-2013 expenditures and significant negative impacts for 2014-2020 expenditures. In less developed regions, the impacts are primarily non-significant, with some negative impacts on the cyclical emission component for 2014-2020 expenditures in certain specifications.
For CH4 emissions, low-carbon expenditures for the 2007-2013 programme show consistently significant negative impacts for all regions, indicating effective reductions on the cyclical emissions component. The 2014-2020 expenditures yield mixed results, with generally non-significant impacts except for significant negative impacts in transition regions. Regarding N2O emissions, low-carbon expenditures for both the 2007-2013 and 2014-2020 programmes typically show non-significant impacts. In transition regions, there are some negative differential impacts for 2014-2020 expenditures, but these are not consistently significant. In less developed regions, the differential impacts are mixed, with positive and significant for the 2007-2013 period and non-significant for 2014-2020.
The estimations for the cyclical emissions components reveal that low-carbon expenditures have mixed results, which vary depending on the type of gas used and the level of regional development. This is observed for both per capita and per unit of GDP.
5 Concluding remarks
This paper examines the impact of climate action through one of the key vehicles for European regions: the “Low-Carbon Economy” funds of the EU Cohesion and Structural Policy for the 2007-2013 and 2014-2020 programmes. Our analysis reveals that the impact of spending on low-carbon economies in reducing GHG emissions varies significantly depending on the level of regional development, according to the EU classification that distinguishes between less developed regions, regions in transition, and developed regions. We observe that, for both analysed programmes, investments in low-carbon economies are associated with a statistically significant decrease in the long-term trend of total emissions, both per capita and relative to GDP, in developed and transition regions. However, in less developed regions, these investments have a statistically significant positive impact, indicating either an increase in emissions or a less effective reduction. Furthermore, estimations based on specific gas types also demonstrate differential impacts.
These results indicate a clear dichotomy: investments yield significant reductions in long-term emissions trends in developed and transition regions but appear to have the opposite effect in less developed regions. This aligns with insights from Doda (2014) and Sheldon (2017), who emphasise the variability of emission responses to economic cycles, particularly noting the stabilising effect of wealthier economies on emission patterns. Conversely, in less developed regions, the counterintuitive increase in emissions suggests a subtle interplay between economic development and environmental sustainability efforts, reminiscent of the disparities identified in the existing literature (Naqvi and Zwickl, 2017; Ivanova et al., 2017).
In addition, our analysis showed that European regions’ responses to low-carbon economies funds vary significantly among programmes, emphasising the importance of taking a nuanced approach when evaluating the impact of these investments on environmental sustainability. This involves considering the particularities of each program, the intrinsic differences between regions and types of emissions.
Our findings support the idea of a refined approach to the implementation of climate policy objectives through the EU cohesion policy. We agree with other studies claiming that a diverse set of guidelines must be implemented to effectively reduce emissions rather than relying on individual initiatives (Brinkley, 2014). Consequently, economy-wide assessments, sector-specific trends, planning expertise in implementing place-based approaches, and extensive case studies can help underscore the complexity of decoupling economic growth from pollutant emissions (Brinkley, 2014; Naqvi and Zwickl, 2017).
In general, the need for differentiated policies to achieve economic development is an issue that has been widely studied since the creation of the Cohesion Policy for European regions, underscoring the importance of addressing persistent inequalities between EU regions. Very early on, Boldrin and Canova (2001) highlighted that these policies tend to be more redistributive than growth-oriented, reflecting political compromises over economic efficiency. However, investments in education and human capital, which constitute a significant portion of these funds, have shown medium-term positive impacts, emphasising the critical role of human capital development in fostering long-term regional growth and convergence (Rodríguez-Pose and Fratesi, 2004). Moreover, Rodríguez-Pose and Garcilazo (2015) and Charron et al. (2014) argue that the quality of local and regional governance directly affects economic growth and the effectiveness of structural funds. They find significant variations in governance quality within countries, such as levels of corruption and efficiency of public services, which have substantial implications for policy design. Iammarino et al. (2019) emphasise the need for tailored policies that are sensitive to the structural prospects of different regions, aligning with the Barca report’s advocacy for place-based strategies (Barca, 2009). These insights align with the broader literature on the Environmental Kuznets Hypothesis, which suggests that economic development can lead to improved environmental outcomes, but only if accompanied by effective governance and tailored policy interventions.
Our research underscores the pressing need for tailored environmental policies to optimise interventions across European regions. The differential impacts of cohesion policy investments on regional emission trends and cycles highlight the necessity of moving away from one-size-fits-all solutions towards development-sensitive and emission-specific strategies. Special attention should be given to less developed regions, which, although not major emitters per capita currently, risk becoming significant pollution producers as they develop economically. Aligning investment strategies with regional profiles can enhance the effectiveness of emissions reduction efforts, contributing to a more nuanced policy dialogue on sustainable development. This underscores the critical intersection of regional growth, sustainability investments, and environmental outcomes within the EU’s policy framework.

### v2

@@ 4.4 Effect by type of gases
| CH4 per capita | CH4 intensity | |||||
| Low Carbon Exp. pc | -0.115 | -0.005 | -0.045 | -0.537*** | -0.054 | -0.147 |
| (0.181) | (0.182) | (0.175) | (0.168) | (0.215) | (0.210) | |
| Low Carbon Exp. pc | 0.826 | 0.042 | -0.095 | -0.109 | -0.275 | -0.573 |
| (0.909) | (0.801) | (0.685) | (1.084) | (0.771) | (0.699) | |
| Low Carbon Exp. pcTER | -0.718** | -0.507 | -0.588 | 0.429 | -0.354 | -0.324 |
| (0.333) | (0.426) | (0.468) | (0.367) | (0.280) | (0.290) | |
| Low Carbon Exp. pcTER | -2.247** | -1.191 | -1.261 | -0.621 | -1.390 | -0.894 |
| (1.136) | (1.103) | (1.103) | (1.215) | (1.108) | (1.087) | |
| Low Carbon Exp. pcLDR | 0.515** | 0.504** | 0.523** | 0.364 | 0.339 | 0.410 |
| (0.201) | (0.199) | (0.211) | (0.238) | (0.271) | (0.275) | |
| Low Carbon Exp. pcLDR | -0.435 | 0.148 | 0.119 | -1.083 | -0.227 | -0.112 |
| (0.887) | (0.777) | (0.679) | (1.040) | (0.764) | (0.700) | |
| N2O per capita | N2O intensity | |||||
| Low Carbon Exp. pc | -0.115 | -0.059 | -0.058 | -0.543*** | -0.092 | -0.163 |
| (0.100) | (0.091) | (0.070) | (0.133) | (0.131) | (0.108) | |
| Low Carbon Exp. pc | 0.180 | 0.142 | 0.288 | -0.774 | -0.118 | -0.196 |
| (0.340) | (0.382) | (0.339) | (0.631) | (0.312) | (0.299) | |
| Low Carbon Exp. pcTER | -0.574*** | -0.187 | -0.140 | 0.570 | -0.088 | 0.134 |
| (0.206) | (0.166) | (0.160) | (0.583) | (0.364) | (0.361) | |
| Low Carbon Exp. pcTER | -0.969*** | -0.764* | -0.829 | 0.647 | -1.049** | -0.442 |
| (0.369) | (0.440) | (0.527) | (0.756) | (0.411) | (0.659) | |
| Low Carbon Exp. pcLDR | 0.536*** | 0.347** | 0.374** | 0.380** | 0.172 | 0.267 |
| (0.178) | (0.144) | (0.150) | (0.193) | (0.191) | (0.176) | |
| Low Carbon Exp. pcLDR | 1.358*** | 0.491 | 0.315 | 0.726 | 0.089 | 0.089 |
| (0.424) | (0.377) | (0.355) | (0.619) | (0.309) | (0.305) | |
| Notes: Estimated coefficients of control variables are omitted. Model specifications correspond to those in Table 2. Robust standard errors clustered at the NUTS 2 level are reported in parentheses. Significance levels: ∗p0.1; ∗∗p0.05; ∗∗∗p0.01. | ||||||
Turning to CH4 emissions (middle panel of Table 6), the 2007–2013 programme does not significantly affect per capita emissions in the control group. There is a negative and significant interaction with transition regions in column 1. In contrast, interactions with LDR are consistently positive and significant. For CH4 intensity, a significant negative impact is observed for the reference group in column 4, but not for the interactions with TER or LDR. Under the 2014–2020 programme, we find no statistically significant effects on either CH4 per capita or intensity, across all regional groups.
Finally, in the case of N2O emissions (lower panel of Table 6), the 2007–2013 programme shows a significant reduction in per capita emissions for TER in column 1, while interaction terms for LDR are again positive and significant throughout. Regarding N2O intensity, a negative and significant coefficient is observed for the reference group in column 4, but positive and significant interactions with LDR offset this. For the 2014–2020 programme, the interaction with TER is negative and statistically significant in both column 1 and 2 (N2O per capita) and column 5 (N2O intensity). The interaction with LDR remains positive and significant only in column 1, in line with the pattern of increased emissions observed in the earlier programme.
As regards the cycle component of CO2 emissions (upper panel of Table 7), the 2007–2013 programme is associated with a statistically significant reduction in per capita emissions for the reference group (developed regions) in models 2 and 3. Additionally, the interaction terms with less developed regions are consistently negative and statistically significant across all specifications. These suggest that low-carbon investments contributed to smoothing short-term fluctuations in CO2 emissions in these areas. The interaction terms with transition regions, however, are not statistically significant. For CO2 intensity, no significant effects are observed for the reference group. However, the interaction terms with LDR are negative and statistically significant in models 5 and 6. For the 2014–2020 programming period, none of the estimated effects are statistically significant for either the reference group or interaction terms.
| Dependent variables by gas type | ||||||
| CO2 per capita | CO2 intensity | |||||
| Model | (1) | (2) | (3) | (4) | (5) | (6) |
| Low Carbon Exp. pc | -0.010 | -0.066* | -0.081** | 0.034 | -0.007 | -0.018 |
| (0.029) | (0.037) | (0.034) | (0.047) | (0.038) | (0.035) | |
| Low Carbon Exp. pc | -0.174 | -0.224 | -0.237 | -0.137 | -0.203 | -0.217 |
| (0.273) | (0.293) | (0.258) | (0.291) | (0.291) | (0.273) | |
| Low Carbon Exp. pcTER | -0.206 | -0.163 | -0.166 | -0.148 | 0.015 | 0.008 |
| (0.165) | (0.145) | (0.143) | (0.197) | (0.113) | (0.118) | |
| Low Carbon Exp. pcTER | 0.325 | 0.090 | 0.127 | -0.113 | -0.478 | -0.459 |
| (0.880) | (0.865) | (0.823) | (0.948) | (0.844) | (0.841) | |
| Low Carbon Exp. pcLDR | -0.152** | -0.169*** | -0.199*** | -0.130* | -0.155*** | -0.173*** |
| (0.065) | (0.062) | (0.061) | (0.068) | (0.054) | (0.051) | |
| Low Carbon Exp. pcLDR | 0.295 | 0.350 | 0.410 | 0.181 | 0.278 | 0.316 |
| (0.285) | (0.302) | (0.274) | (0.298) | (0.291) | (0.275) | |
| CH4 per capita | CH4 intensity | |||||
| Low Carbon Exp. pc | -0.085*** | -0.073*** | -0.062*** | -0.041 | -0.014 | 0.000 |
| (0.025) | (0.023) | (0.022) | (0.027) | (0.028) | (0.031) | |
| Low Carbon Exp. pc | 0.252** | 0.148 | 0.145 | 0.289*** | 0.169* | 0.165* |
| (0.112) | (0.103) | (0.111) | (0.106) | (0.092) | (0.089) | |
| Low Carbon Exp. pcTER | -0.109 | -0.062 | -0.035 | -0.050 | 0.117** | 0.140*** |
| (0.094) | (0.065) | (0.058) | (0.127) | (0.047) | (0.053) | |
| Low Carbon Exp. pcTER | -0.527 | -0.614* | -0.667** | -0.965** | -1.182*** | -1.252*** |
| (0.354) | (0.337) | (0.334) | (0.411) | (0.304) | (0.337) | |
| Low Carbon Exp. pcLDR | 0.019 | 0.000 | 0.010 | 0.040 | 0.014 | 0.036 |
| (0.051) | (0.055) | (0.055) | (0.054) | (0.058) | (0.058) | |
| Low Carbon Exp. pcLDR | -0.256 | 0.037 | -0.002 | -0.369** | -0.035 | -0.097 |
| (0.172) | (0.173) | (0.184) | (0.170) | (0.178) | (0.177) | |
| N2O per capita | N2O intensity | |||||
| Low Carbon Exp. pc | -0.009 | -0.022 | -0.028 | 0.036 | 0.038 | 0.035 |
| (0.021) | (0.029) | (0.031) | (0.026) | (0.025) | (0.028) | |
| Low Carbon Exp. pc | 0.093 | 0.066 | 0.048 | 0.137 | 0.093 | 0.074 |
| (0.150) | (0.149) | (0.127) | (0.166) | (0.146) | (0.142) | |
| Low Carbon Exp. pcTER | -0.098** | -0.159*** | -0.163*** | -0.040 | 0.020 | 0.011 |
| (0.044) | (0.056) | (0.057) | (0.069) | (0.039) | (0.046) | |
| Low Carbon Exp. pcTER | 0.050 | 0.015 | 0.017 | -0.381 | -0.549*** | -0.563*** |
| (0.259) | (0.268) | (0.253) | (0.245) | (0.198) | (0.212) | |
| Low Carbon Exp. pcLDR | 0.113* | 0.129* | 0.121* | 0.134** | 0.143* | 0.147** |
| (0.063) | (0.073) | (0.072) | (0.064) | (0.073) | (0.073) | |
| Low Carbon Exp. pcLDR | -0.167 | -0.083 | -0.068 | -0.284* | -0.159 | -0.166 |
| (0.153) | (0.162) | (0.148) | (0.165) | (0.158) | (0.156) | |
| Notes: Estimated coefficients of control variables are omitted. Model specifications correspond to those in Table 2. Robust standard errors clustered at the NUTS 2 level are reported in parentheses. Significance levels: ∗p0.1; ∗∗p0.05; ∗∗∗p0.01. | ||||||
Turning to the cycle component of CH4 emissions (middle panel of Table 7), the 2007–2013 programme is associated with consistent and statistically significant reductions in per capita emissions for the reference group across all model specifications. In contrast, interaction terms with both TER and LDR are not statistically significant. In the case of the 2014–2020 programme, we observe a positive and statistically significant effect for the reference group in model 1 (CH4 per capita), and across all models for CH4 intensity (columns 4–6). The interaction terms with TER are negative and statistically significant in all models, except model 1. Interaction terms with LDR are not significant in most models, except CH4 intensity in column 4, where a negative effect is detected.
Regarding the cycle component of N2O emissions (lower panel of Table 7), the 2007–2013 programme shows no significant effects for the reference group in terms of per capita emissions. However, all estimated interaction terms with TER are negative and statistically significant, while those with LDR are positive and significant (columns 1-3). For N2O intensity, the reference group shows no statistically significant response. However, the interaction terms with LDR are positive and statistically significant in models 4–6. For the 2014–2020 programming period, interaction terms with TER are again negative and statistically significant in models 5 and 6, and for LDR in model 1.
5 Discussion
The results from the baseline fixed effects model reveal heterogeneous impacts of low-carbon investments on both emission trends and cycles. Generally, these investments do not yield statistically significant outcomes in more developed regions. However, in less developed areas, low-carbon investments appear paradoxically associated with increases in the long-term trend of per capita emissions, and, in some specifications, also in emissions intensity. This suggests that, in these regions, investments stimulate economic growth that remains dependent on carbon-intensive technologies. In contrast, when examining short-term fluctuations, the results suggest a more positive role for these investments in reducing emissions volatility, especially in less developed regions, indicating that the funds may help stabilise emissions in the short run.
When we employ instrumental variables to address potential endogeneity, the findings further reinforce the existence of regional disparities. The positive effect on per capita emissions in less developed regions is clearly confirmed, and additional significant effects emerge for transition regions that were not visible in the initial analysis. While this might suggest that estimations for transition regions were underestimated, the overall pattern is strengthened, reflecting heterogeneous regional impacts. Hence, the IV analysis enhances the credibility of a causal interpretation of the differentiated expenditure effects.
When disaggregating the results by GHG type (CO2, CH4, and N2O), a more nuanced picture emerges. For trends, the increase in per capita emissions observed in less developed regions is primarily driven by CO2, the gas most closely linked to industrial activity and energy production. In contrast, the results for CH4 and N2O show mixed patterns, with adverse effects in some less developed regions and mitigating effects in transition areas. Regarding emission cycles, low-carbon investments show more consistently positive effects, particularly in reducing volatility for CO2 in less developed regions and for CH4 and N2O in transition regions.
However, these results should be interpreted in light of the characteristics of the emission sources and policy contexts associated with each gas. Unlike CO2, which is closely tied to fossil fuel combustion and therefore directly targeted by ESI-funded low-carbon investments, CH4 and N2O emissions are largely linked to agricultural practices and specific industrial processes (Scheffler and Wiegmann, 2024). Their responsiveness to low-carbon funding may thus be more limited or dependent on complementary policies. In particular, the Common Agricultural Policy (CAP), which is funded separately from the ESI funds, has distinct objectives that may create overlaps with climate-related goals (European Court of Auditors, 2021).
More broadly, the low-carbon economy fund from the ESI funds is one among several instruments in the EU’s climate policy architecture. Notably, the EU ETS has played a more central role in recent decades (Bayer and Aklin, 2020). Nevertheless, ESI funding remains important, as it aims to address regional disparities and supports place-based transitions.
Altogether, our findings provide strong empirical support for our central hypothesis. While we find consistent evidence that low-carbon economy funding does not lead to long-term reductions in emissions indicators in more developed regions, the adverse effects observed in less developed regions are particularly concerning, given that these areas are the primary targets of ESI fund support. Nevertheless, the results suggest that while the funds may not be highly effective in driving long-term decarbonisation, they could still play a stabilising role by dampening short-term emissions volatility.
It is important to note that most statistically significant effects are concentrated in the 2007-2013 programming period. The lack of robust results for the 2014-2020 period may be due to data limitations. Given that the environmental impacts of investments often take time to materialise, the imprecision likely reflects the short post-investment observation window.
Moreover, in models using emissions intensity as the dependent variable, results are more frequently statistically not significant. Although this did not invalidate the central hypothesis of the study, it raises policy concerns. Specifically, the limited effects on emissions intensity, even in more developed areas, clearly hint that funding is not effectively supporting the decoupling of emissions from economic growth.
These findings call for a reassessment of EU climate investment strategies. On the one hand, the failure to curb long-term emissions in less developed areas raises questions about allocative efficiency and the risk of reinforcing carbon lock-in (Unruh, 2000). On the other hand, mechanisms behind the heterogeneous impacts – such as institutional capacity gaps, infrastructure rigidity, or lack of policy complementarities – remain underexplored and merit further attention (Rodríguez-Pose and Garcilazo, 2015; Charron et al., 2014). In this respect, Iammarino et al. (2019) emphasise the need for tailored policies that are sensitive to the structural prospects of different regions, aligning with the Barca report’s advocacy for place-based strategies (Barca, 2009).
In addition, while the observed short-term stabilisation of emissions is encouraging, it may reflect counter-cyclical spending effects rather than deeper structural transformations, thereby questioning the long-term decarbonisation potential of these funds.
Finally, although beyond this paper’s scope, some of the emissions increases in less developed regions may reflect industrial relocation rather than policy failure. From a broader EU or global perspective, such shifts may be environmentally neutral. These trade-offs also underline the need for more targeted, region-specific strategies. Future research should investigate the mechanisms behind regional disparities and identify the complementary policies needed to align the ESI spending with long-term climate and development objectives.
6 Conclusions
This paper examines the climate impact of one of the European Union’s key instruments for regional development: the “Low-Carbon Economy” funds allocated through the European Structural and Investment funds during the 2007-2013 and 2014-2020 programming periods. Focusing on per capita emissions and emissions intensity – both decomposed into trend and cyclical components – we estimate the structural (long-run) and cyclical (short-run) effects of low-carbon expenditures across regions.
Our findings point to a substantial degree of heterogeneity. In more developed and transition regions, low-carbon funding shows no consistent long-term effect in reducing emissions, whether measured per capita or relative to GDP. However, such investments are associated with significant increases in long-term emissions trends in less developed regions, despite being the primary beneficiaries of these funds. These results suggest that, in lagging regions, climate-related investments may stimulate economic activity without achieving the intended decoupling from carbon-intensive processes. Instrumental variable estimations reinforce the robustness of our findings and lend support to a causal interpretation of the observed relationships.
When disaggregating emissions by gas type (CO2, CH4, and N2O), results show further heterogeneity. In less developed regions, increases in emissions are mainly associated with CO2, whereas the effects on methane and nitrous oxide are less robust and vary across regions. Regarding short-run dynamics, we find some evidence that low-carbon investments contributed to reducing emissions volatility, especially related to expenditures of the 2007-2013 programme.
Overall, the empirical results raise questions about how effectively the EU Cohesion Policy supports the broader decarbonisation agenda. While our findings point to some limitations in the environmental impact of low-carbon investments, they also highlight the regional challenges in turning funding into measurable emission reductions. Structural conditions, institutional capacity, and how policies are implemented matter. However, other elements, such as the types of projects financed, their interaction with other policy initiatives, and the time it takes for their effects to materialise, likely play a role as well.
Our findings suggest that improving the climate effectiveness of the Cohesion Policy will require a more substantial alignment with regional contexts and development trajectories. Given the heterogeneous impacts observed across regions, it is essential to move beyond uniform approaches. Less developed regions deserve particular attention. Although they are not currently the highest emitters on a per capita basis, they face the risk of becoming major sources of emissions as they develop. We claim that aligning investment strategies with regional profiles can enhance the effectiveness of emissions reduction efforts.